OpenStudy (arihangdu):

The equation C=5/9(F-32) relates temperatures in Fahrenheit F and degrees Celsius C. What is F in terms of C?

OpenStudy (michele_laino):

If I multiply both sides of your formula by \(9/5\), I get: \[C \cdot \frac{9}{5} = \frac{5}{9}\left( {F - 32} \right) \cdot \frac{9}{5}\] please simplify

OpenStudy (arihangdu):

Then?

OpenStudy (michele_laino):

after the simplification, I get: \[\frac{9}{5}C = F - 32\] then I add 32 to both sides, so I can write this: \[\frac{9}{5}C + 32 = F - 32 + 32\] please try to simplify

OpenStudy (michele_laino):

at the right side we have: \(-32+32=0\), so, what do you get?

OpenStudy (arihangdu):

Is that F = 9/5C+32 ?

OpenStudy (michele_laino):

correct!

OpenStudy (arihangdu):

Really ?

OpenStudy (michele_laino):

yes! It is correct

OpenStudy (arihangdu):

Thanks

OpenStudy (michele_laino):

:)